# NOLO: Navigate Only Look Once

### v1

@@ D.1 RoboTHOR Video Navigation Performance
FloorPlan_Train10_5 72.94$\pm$ 5.08 27.83$\pm$ 0.15 252.83$\pm$ 8.62 1.86$\pm$ 0.29
FloorPlan_Train11_5 82.00$\pm$ 1.44 36.28$\pm$ 2.90 195.77$\pm$ 8.38 1.50$\pm$ 0.05
FloorPlan_Train12_5 73.53$\pm$ 0.83 32.13$\pm$ 1.26 230.53$\pm$ 8.63 1.59$\pm$ 0.05
FloorPlan_Train1_5 64.72$\pm$ 1.04 32.41$\pm$ 1.39 258.61$\pm$ 2.70 1.94$\pm$ 0.06
FloorPlan_Train2_5 70.63$\pm$ 3.40 29.21$\pm$ 3.18 243.38$\pm$ 16.56 1.78$\pm$ 0.11
FloorPlan_Train3_5 62.16$\pm$ 1.94 23.55$\pm$ 0.62 288.69$\pm$ 8.45 2.27$\pm$ 0.09
FloorPlan_Train4_5 72.12$\pm$ 5.77 27.65$\pm$ 0.21 244.76$\pm$ 16.37 1.89$\pm$ 0.20
FloorPlan_Train5_5 75.61$\pm$ 0.57 32.77$\pm$ 1.68 232.83$\pm$ 3.66 1.60$\pm$ 0.04
FloorPlan_Train6_5 67.78$\pm$ 4.37 27.60$\pm$ 0.31 254.15$\pm$ 5.21 2.00$\pm$ 0.12
FloorPlan_Train7_5 75.62$\pm$ 1.84 19.90$\pm$ 2.02 256.17$\pm$ 8.42 1.55$\pm$ 0.03
FloorPlan_Train8_5 65.62$\pm$ 2.55 26.02$\pm$ 2.55 253.82$\pm$ 8.94 2.00$\pm$ 0.07
FloorPlan_Train9_5 56.67$\pm$ 4.11 24.40$\pm$ 2.52 303.16$\pm$ 8.15 2.18$\pm$ 0.05
@@
FloorPlan_Val1_1 64.33$\pm$ 5.44 25.29$\pm$ 5.10 283.93$\pm$ 29.92 2.12$\pm$ 0.22
FloorPlan_Val1_2 57.42$\pm$ 3.57 22.97$\pm$ 1.36 305.57$\pm$ 10.25 2.34$\pm$ 0.15
FloorPlan_Val1_3 82.98$\pm$ 1.38 41.56$\pm$ 1.40 197.25$\pm$ 10.54 1.43$\pm$ 0.04
FloorPlan_Val1_4 62.22$\pm$ 3.27 25.73$\pm$ 1.58 284.70$\pm$ 10.51 2.26$\pm$ 0.10
FloorPlan_Val1_5 58.61$\pm$ 1.42 31.95$\pm$ 2.50 276.70$\pm$ 4.67 2.53$\pm$ 0.12
FloorPlan_Val2_1 52.73$\pm$ 1.96 17.41$\pm$ 1.69 336.15$\pm$ 1.47 2.59$\pm$ 0.08
FloorPlan_Val2_2 67.22$\pm$ 4.33 28.52$\pm$ 2.17 268.21$\pm$ 2.39 1.95$\pm$ 0.07
FloorPlan_Val2_3 79.74$\pm$ 3.46 35.81$\pm$ 0.85 209.43$\pm$ 23.21 1.34$\pm$ 0.08
FloorPlan_Val2_4 62.95$\pm$ 6.01 25.65$\pm$ 0.77 276.85$\pm$ 16.89 2.10$\pm$ 0.19
FloorPlan_Val2_5 58.70$\pm$ 0.94 24.36$\pm$ 1.83 290.25$\pm$ 2.98 2.55$\pm$ 0.05
FloorPlan_Val3_1 71.43$\pm$ 2.33 26.32$\pm$ 3.77 240.97$\pm$ 20.02 2.00$\pm$ 0.14
FloorPlan_Val3_2 71.53$\pm$ 1.75 26.78$\pm$ 0.32 238.58$\pm$ 3.53 1.99$\pm$ 0.09
FloorPlan_Val3_3 72.59$\pm$ 4.57 27.82$\pm$ 1.74 242.52$\pm$ 9.40 1.73$\pm$ 0.18
FloorPlan_Val3_4 57.30$\pm$ 2.65 18.35$\pm$ 0.51 300.88$\pm$ 11.16 2.30$\pm$ 0.08
FloorPlan_Val3_5 70.97$\pm$ 3.16 29.52$\pm$ 1.81 243.77$\pm$ 19.50 1.96$\pm$ 0.14
@@
Table 9: NOLO-T Performance on RoboTHOR Testing Set
@@
FloorPlan_Train10_5 76.47$\pm$ 3.53 29.97$\pm$ 2.31 225.76$\pm$ 18.05 1.71$\pm$ 0.16
FloorPlan_Train11_5 78.00$\pm$ 0.67 36.00$\pm$ 0.43 210.91$\pm$ 7.94 1.56$\pm$ 0.10
FloorPlan_Train12_5 75.88$\pm$ 0.59 30.77$\pm$ 1.29 233.25$\pm$ 4.82 1.62$\pm$ 0.02
FloorPlan_Train1_5 64.17$\pm$ 2.50 32.47$\pm$ 4.30 254.97$\pm$ 6.37 2.07$\pm$ 0.01
FloorPlan_Train2_5 65.71$\pm$ 6.67 26.12$\pm$ 4.84 259.38$\pm$ 29.69 1.95$\pm$ 0.25
FloorPlan_Train3_5 59.41$\pm$ 4.12 22.67$\pm$ 0.98 303.37$\pm$ 11.79 2.39$\pm$ 0.08
FloorPlan_Train4_5 72.73$\pm$ 0.91 27.02$\pm$ 2.10 238.05$\pm$ 10.76 1.89$\pm$ 0.09
FloorPlan_Train5_5 78.18$\pm$ 3.64 33.88$\pm$ 0.94 224.37$\pm$ 22.06 1.48$\pm$ 0.12
FloorPlan_Train6_5 74.17$\pm$ 0.83 32.28$\pm$ 1.91 216.92$\pm$ 17.75 1.83$\pm$ 0.11
FloorPlan_Train7_5 80.62$\pm$ 0.62 23.85$\pm$ 1.03 232.86$\pm$ 15.75 1.38$\pm$ 0.08
FloorPlan_Train8_5 66.25$\pm$ 2.50 30.95$\pm$ 2.62 245.14$\pm$ 15.49 2.06$\pm$ 0.12
FloorPlan_Train9_5 61.00$\pm$ 1.00 28.92$\pm$ 1.60 285.27$\pm$ 2.75 1.80$\pm$ 0.20
@@
FloorPlan_Val1_1 64.00$\pm$ 2.00 26.54$\pm$ 1.03 277.16$\pm$ 12.08 2.25$\pm$ 0.02
FloorPlan_Val1_2 64.55 25.30$\pm$ 0.81 272.34$\pm$ 12.70 2.20$\pm$ 0.01
FloorPlan_Val1_3 80.53$\pm$ 5.79 36.93$\pm$ 0.98 201.32$\pm$ 9.73 1.62$\pm$ 0.22
FloorPlan_Val1_4 59.44$\pm$ 7.22 25.50$\pm$ 1.25 283.97$\pm$ 24.48 2.14$\pm$ 0.21
FloorPlan_Val1_5 62.50$\pm$ 0.83 32.49$\pm$ 0.23 268.84$\pm$ 5.32 2.06$\pm$ 0.01
FloorPlan_Val2_1 56.36$\pm$ 0.91 17.59$\pm$ 0.30 329.25$\pm$ 10.75 2.44$\pm$ 0.04
FloorPlan_Val2_2 65.56$\pm$ 2.22 30.85$\pm$ 2.09 256.13$\pm$ 9.46 2.07$\pm$ 0.07
FloorPlan_Val2_3 76.92$\pm$ 1.54 34.22$\pm$ 0.03 209.07$\pm$ 4.52 1.59$\pm$ 0.02
FloorPlan_Val2_4 68.08$\pm$ 0.38 25.24$\pm$ 0.19 265.69$\pm$ 6.21 1.93$\pm$ 0.00
FloorPlan_Val2_5 65.22$\pm$ 3.48 24.79$\pm$ 0.74 290.28$\pm$ 11.04 2.29$\pm$ 0.14
FloorPlan_Val3_1 69.52$\pm$ 0.95 25.09$\pm$ 0.17 244.60$\pm$ 10.47 2.04$\pm$ 0.03
FloorPlan_Val3_2 75.83$\pm$ 0.83 27.14$\pm$ 2.16 223.59$\pm$ 6.15 1.81$\pm$ 0.06
FloorPlan_Val3_3 73.70$\pm$ 2.59 24.63$\pm$ 2.77 245.72$\pm$ 15.48 1.75$\pm$ 0.15
FloorPlan_Val3_4 52.38$\pm$ 3.81 17.61$\pm$ 0.82 316.57$\pm$ 4.58 2.46$\pm$ 0.12
FloorPlan_Val3_5 76.67$\pm$ 0.83 31.29$\pm$ 1.91 211.91$\pm$ 5.29 1.62$\pm$ 0.04
@@
Table 10: NOLO(M) Performance on RoboTHOR Testing Set
@@
FloorPlan_Train10_5 80.59$\pm$ 4.12 32.92$\pm$ 1.93 202.68$\pm$ 22.39 1.47$\pm$ 0.12
FloorPlan_Train11_5 77.33$\pm$ 2.67 36.00$\pm$ 3.27 218.04$\pm$ 1.64 1.64$\pm$ 0.01
FloorPlan_Train12_5 78.82 36.62$\pm$ 0.82 211.44$\pm$ 9.89 1.55$\pm$ 0.01
FloorPlan_Train1_5 50.83$\pm$ 5.83 27.99$\pm$ 2.85 302.14$\pm$ 14.69 2.35$\pm$ 0.10
FloorPlan_Train2_5 69.05$\pm$ 1.43 25.09$\pm$ 2.83 256.49$\pm$ 1.55 1.85$\pm$ 0.13
FloorPlan_Train3_5 54.71$\pm$ 0.59 21.49$\pm$ 0.94 312.69$\pm$ 3.18 2.61$\pm$ 0.08
FloorPlan_Train4_5 73.64$\pm$ 2.73 25.06$\pm$ 1.03 245.16$\pm$ 17.39 1.78$\pm$ 0.08
FloorPlan_Train5_5 75.45$\pm$ 0.91 31.23$\pm$ 1.14 234.16$\pm$ 1.94 1.75$\pm$ 0.07
FloorPlan_Train6_5 65.83$\pm$ 2.50 24.81$\pm$ 0.36 276.04$\pm$ 19.72 1.95$\pm$ 0.01
FloorPlan_Train7_5 70.62$\pm$ 0.63 19.64$\pm$ 0.81 281.29$\pm$ 2.88 1.77$\pm$ 0.05
FloorPlan_Train8_5 59.38$\pm$ 6.87 23.84$\pm$ 2.27 287.47$\pm$ 17.78 2.38$\pm$ 0.37
FloorPlan_Train9_5 59.00$\pm$ 7.00 31.86$\pm$ 5.95 278.41$\pm$ 28.49 2.12$\pm$ 0.22
@@
FloorPlan_Val1_1 60.00$\pm$ 3.00 25.32$\pm$ 3.23 294.85$\pm$ 21.57 2.29$\pm$ 0.09
FloorPlan_Val1_2 60.00$\pm$ 1.82 24.32$\pm$ 3.48 291.72$\pm$ 6.24 2.18$\pm$ 0.09
FloorPlan_Val1_3 82.63$\pm$ 1.58 35.75$\pm$ 3.13 200.15$\pm$ 3.24 1.41$\pm$ 0.09
FloorPlan_Val1_4 62.22$\pm$ 1.11 24.95$\pm$ 0.16 288.34$\pm$ 6.13 2.19$\pm$ 0.21
FloorPlan_Val1_5 64.58$\pm$ 4.58 31.77$\pm$ 3.05 265.98$\pm$ 16.41 2.04$\pm$ 0.16
FloorPlan_Val2_1 49.09$\pm$ 1.82 14.51$\pm$ 2.21 358.64$\pm$ 2.52 2.68$\pm$ 0.06
FloorPlan_Val2_2 66.11$\pm$ 2.78 30.06$\pm$ 1.07 261.83$\pm$ 10.29 2.11$\pm$ 0.15
FloorPlan_Val2_3 77.69$\pm$ 0.77 31.33$\pm$ 1.61 232.93$\pm$ 6.27 1.49$\pm$ 0.02
FloorPlan_Val2_4 57.69$\pm$ 1.54 23.61$\pm$ 0.69 296.56$\pm$ 8.78 2.26$\pm$ 0.07
FloorPlan_Val2_5 59.57$\pm$ 4.78 23.29$\pm$ 0.57 303.90$\pm$ 19.02 2.48$\pm$ 0.30
FloorPlan_Val3_1 70.48$\pm$ 1.90 25.14$\pm$ 1.24 254.00$\pm$ 0.10 1.87$\pm$ 0.06
FloorPlan_Val3_2 71.25$\pm$ 0.42 25.59$\pm$ 0.67 244.68$\pm$ 2.23 1.88$\pm$ 0.04
FloorPlan_Val3_3 67.78$\pm$ 0.37 24.20$\pm$ 1.78 258.00$\pm$ 2.40 1.89$\pm$ 0.03
FloorPlan_Val3_4 54.29$\pm$ 0.95 13.15$\pm$ 1.33 330.42$\pm$ 8.02 2.40$\pm$ 0.00
FloorPlan_Val3_5 67.50$\pm$ 1.67 30.21$\pm$ 4.40 246.60$\pm$ 8.73 2.05$\pm$ 0.04
@@
D.2 Habitat Video Navigation Performance
Table 11: NOLO Performance on Habitat Testing Set

### v2

@@ 9.1 RoboTHOR Evaluation Performance
T10_5 33.84 9.29 422.72 1.92
T11_5 36.72 11.13 434.58 2.22
T12_5 31.24 8.14 437.61 2.12
T1_5 28.98 7.44 428.73 2.00
T2_5 32.59 9.05 440.12 2.14
T3_5 24.53 8.26 444.36 2.45
T4_5 35.21 10.57 423.94 1.99
T5_5 33.47 10.36 433.51 2.09
T6_5 39.12 11.98 415.89 2.32
T7_5 25.74 6.92 439.82 2.18
T8_5 36.23 11.45 421.16 2.27
T9_5 32.67 8.54 426.75 2.15
Average 32.53 9.43 430.77 2.15
V1_1 27.14 8.21 462.48 2.65
V1_2 29.62 11.18 440.87 2.53
V1_3 30.28 15.24 405.27 1.84
V1_4 22.84 7.91 454.28 2.59
V1_5 30.41 6.76 447.91 2.21
V2_1 34.32 13.43 425.68 2.05
V2_2 48.45 10.92 398.67 1.25
V2_3 43.67 12.57 387.45 1.73
V2_4 36.18 14.97 404.22 1.49
V2_5 31.84 12.98 412.88 2.29
V3_1 11.14 5.23 450.75 3.07
V3_2 28.89 19.34 409.38 1.97
V3_3 34.67 13.15 429.52 2.21
V3_4 27.93 10.27 439.74 2.55
V3_5 29.78 11.33 399.28 1.53
Average 31.14 11.57 424.56 2.13
Table 10: VGM performance on RoboTHOR testing set. T” denotes FloorPlan_Train and V” denotes FloorPlan_Val.
[…]
T10_5 46.02 25.47 375.88 2.49
T11_5 44.98 23.65 392.45 2.71
T12_5 50.13 22.71 378.29 2.39
T1_5 49.75 26.02 371.15 2.55
T2_5 43.59 21.54 388.67 2.23
T3_5 46.93 22.88 367.36 2.67
T4_5 48.32 24.07 381.74 2.41
T5_5 44.25 24.61 373.28 2.44
T6_5 50.41 25.19 369.24 2.33
T7_5 45.67 22.32 365.72 2.62
T8_5 46.89 23.78 380.91 2.35
T9_5 47.54 24.56 378.32 2.13
Average 47.04 23.90 376.92 2.44
V1_1 41.13 21.57 405.87 2.64
V1_2 39.23 19.31 412.88 2.87
V1_3 43.67 22.13 386.24 1.77
V1_4 42.45 18.27 416.54 2.41
V1_5 35.88 18.68 401.96 2.54
V2_1 44.26 20.67 382.33 2.38
V2_2 49.87 23.12 394.56 1.85
V2_3 41.22 21.87 385.19 1.91
V2_4 38.96 22.07 396.72 2.25
V2_5 43.78 20.52 405.33 2.64
V3_1 35.37 12.43 420.64 3.22
V3_2 39.16 21.38 389.33 2.42
V3_3 33.78 18.04 406.57 2.11
V3_4 30.15 14.92 416.38 2.82
V3_5 48.76 24.87 392.94 1.87
Average 40.51 19.99 400.90 2.38
Table 11: ZSON performance on RoboTHOR testing set. T” denotes FloorPlan_Train and V” denotes FloorPlan_Val.
[…]
T10_5 38.89 20.69 350.67 2.23
T11_5 35.06 19.92 362.44 2.41
T12_5 40.06 22.54 345.23 2.16
T1_5 36.17 19.45 358.91 2.09
T2_5 34.37 22.58 351.12 2.02
T3_5 41.86 23.24 340.11 2.34
T4_5 38.72 21.52 346.72 2.29
T5_5 36.06 20.35 349.33 2.25
T6_5 43.17 22.61 335.84 2.38
T7_5 35.17 20.45 360.17 2.15
T8_5 35.96 20.27 348.52 2.31
T9_5 38.52 22.08 342.66 2.18
Average 37.83 21.31 349.31 2.23
V1_1 37.91 18.12 378.87 2.72
V1_2 34.50 17.64 386.19 2.91
V1_3 43.06 20.88 359.72 1.77
V1_4 28.71 16.29 395.18 2.56
V1_5 33.77 15.17 380.14 2.64
V2_1 40.93 19.42 361.48 2.12
V2_2 41.60 25.01 353.26 1.43
V2_3 43.34 21.41 368.11 1.85
V2_4 39.55 22.75 354.24 2.25
V2_5 40.02 19.83 360.73 2.72
V3_1 19.52 10.81 392.54 3.22
V3_2 37.89 19.64 368.18 2.56
V3_3 32.41 15.28 382.45 2.21
V3_4 23.77 13.04 390.76 2.89
V3_5 39.79 26.81 358.32 1.73
Average 35.78 18.81 372.68 2.37
Table 12: NOLO-C performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 72.94$\pm$ 5.08 27.83$\pm$ 0.15 252.83$\pm$ 8.62 1.86$\pm$ 0.29
T11_5 82.00$\pm$ 1.44 36.28$\pm$ 2.90 195.77$\pm$ 8.38 1.50$\pm$ 0.05
T12_5 73.53$\pm$ 0.83 32.13$\pm$ 1.26 230.53$\pm$ 8.63 1.59$\pm$ 0.05
T1_5 64.72$\pm$ 1.04 32.41$\pm$ 1.39 258.61$\pm$ 2.70 1.94$\pm$ 0.06
T2_5 70.63$\pm$ 3.40 29.21$\pm$ 3.18 243.38$\pm$ 16.56 1.78$\pm$ 0.11
T3_5 62.16$\pm$ 1.94 23.55$\pm$ 0.62 288.69$\pm$ 8.45 2.27$\pm$ 0.09
T4_5 72.12$\pm$ 5.77 27.65$\pm$ 0.21 244.76$\pm$ 16.37 1.89$\pm$ 0.20
T5_5 75.61$\pm$ 0.57 32.77$\pm$ 1.68 232.83$\pm$ 3.66 1.60$\pm$ 0.04
T6_5 67.78$\pm$ 4.37 27.60$\pm$ 0.31 254.15$\pm$ 5.21 2.00$\pm$ 0.12
T7_5 75.62$\pm$ 1.84 19.90$\pm$ 2.02 256.17$\pm$ 8.42 1.55$\pm$ 0.03
T8_5 65.62$\pm$ 2.55 26.02$\pm$ 2.55 253.82$\pm$ 8.94 2.00$\pm$ 0.07
T9_5 56.67$\pm$ 4.11 24.40$\pm$ 2.52 303.16$\pm$ 8.15 2.18$\pm$ 0.05
@@
V1_1 64.33$\pm$ 5.44 25.29$\pm$ 5.10 283.93$\pm$ 29.92 2.12$\pm$ 0.22
V1_2 57.42$\pm$ 3.57 22.97$\pm$ 1.36 305.57$\pm$ 10.25 2.34$\pm$ 0.15
V1_3 82.98$\pm$ 1.38 41.56$\pm$ 1.40 197.25$\pm$ 10.54 1.43$\pm$ 0.04
V1_4 62.22$\pm$ 3.27 25.73$\pm$ 1.58 284.70$\pm$ 10.51 2.26$\pm$ 0.10
V1_5 58.61$\pm$ 1.42 31.95$\pm$ 2.50 276.70$\pm$ 4.67 2.53$\pm$ 0.12
V2_1 52.73$\pm$ 1.96 17.41$\pm$ 1.69 336.15$\pm$ 1.47 2.59$\pm$ 0.08
V2_2 67.22$\pm$ 4.33 28.52$\pm$ 2.17 268.21$\pm$ 2.39 1.95$\pm$ 0.07
V2_3 79.74$\pm$ 3.46 35.81$\pm$ 0.85 209.43$\pm$ 23.21 1.34$\pm$ 0.08
V2_4 62.95$\pm$ 6.01 25.65$\pm$ 0.77 276.85$\pm$ 16.89 2.10$\pm$ 0.19
V2_5 58.70$\pm$ 0.94 24.36$\pm$ 1.83 290.25$\pm$ 2.98 2.55$\pm$ 0.05
V3_1 71.43$\pm$ 2.33 26.32$\pm$ 3.77 240.97$\pm$ 20.02 2.00$\pm$ 0.14
V3_2 71.53$\pm$ 1.75 26.78$\pm$ 0.32 238.58$\pm$ 3.53 1.99$\pm$ 0.09
V3_3 72.59$\pm$ 4.57 27.82$\pm$ 1.74 242.52$\pm$ 9.40 1.73$\pm$ 0.18
V3_4 57.30$\pm$ 2.65 18.35$\pm$ 0.51 300.88$\pm$ 11.16 2.30$\pm$ 0.08
V3_5 70.97$\pm$ 3.16 29.52$\pm$ 1.81 243.77$\pm$ 19.50 1.96$\pm$ 0.14
@@
Table 13: NOLO-T performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
@@
T10_5 76.47$\pm$ 3.53 29.97$\pm$ 2.31 225.76$\pm$ 18.05 1.71$\pm$ 0.16
T11_5 78.00$\pm$ 0.67 36.00$\pm$ 0.43 210.91$\pm$ 7.94 1.56$\pm$ 0.10
T12_5 75.88$\pm$ 0.59 30.77$\pm$ 1.29 233.25$\pm$ 4.82 1.62$\pm$ 0.02
T1_5 64.17$\pm$ 2.50 32.47$\pm$ 4.30 254.97$\pm$ 6.37 2.07$\pm$ 0.01
T2_5 65.71$\pm$ 6.67 26.12$\pm$ 4.84 259.38$\pm$ 29.69 1.95$\pm$ 0.25
T3_5 59.41$\pm$ 4.12 22.67$\pm$ 0.98 303.37$\pm$ 11.79 2.39$\pm$ 0.08
T4_5 72.73$\pm$ 0.91 27.02$\pm$ 2.10 238.05$\pm$ 10.76 1.89$\pm$ 0.09
T5_5 78.18$\pm$ 3.64 33.88$\pm$ 0.94 224.37$\pm$ 22.06 1.48$\pm$ 0.12
T6_5 74.17$\pm$ 0.83 32.28$\pm$ 1.91 216.92$\pm$ 17.75 1.83$\pm$ 0.11
T7_5 80.62$\pm$ 0.62 23.85$\pm$ 1.03 232.86$\pm$ 15.75 1.38$\pm$ 0.08
T8_5 66.25$\pm$ 2.50 30.95$\pm$ 2.62 245.14$\pm$ 15.49 2.06$\pm$ 0.12
T9_5 61.00$\pm$ 1.00 28.92$\pm$ 1.60 285.27$\pm$ 2.75 1.80$\pm$ 0.20
@@
V1_1 64.00$\pm$ 2.00 26.54$\pm$ 1.03 277.16$\pm$ 12.08 2.25$\pm$ 0.02
V1_2 64.55 25.30$\pm$ 0.81 272.34$\pm$ 12.70 2.20$\pm$ 0.01
V1_3 80.53$\pm$ 5.79 36.93$\pm$ 0.98 201.32$\pm$ 9.73 1.62$\pm$ 0.22
V1_4 59.44$\pm$ 7.22 25.50$\pm$ 1.25 283.97$\pm$ 24.48 2.14$\pm$ 0.21
V1_5 62.50$\pm$ 0.83 32.49$\pm$ 0.23 268.84$\pm$ 5.32 2.06$\pm$ 0.01
V2_1 56.36$\pm$ 0.91 17.59$\pm$ 0.30 329.25$\pm$ 10.75 2.44$\pm$ 0.04
V2_2 65.56$\pm$ 2.22 30.85$\pm$ 2.09 256.13$\pm$ 9.46 2.07$\pm$ 0.07
V2_3 76.92$\pm$ 1.54 34.22$\pm$ 0.03 209.07$\pm$ 4.52 1.59$\pm$ 0.02
V2_4 68.08$\pm$ 0.38 25.24$\pm$ 0.19 265.69$\pm$ 6.21 1.93$\pm$ 0.00
V2_5 65.22$\pm$ 3.48 24.79$\pm$ 0.74 290.28$\pm$ 11.04 2.29$\pm$ 0.14
V3_1 69.52$\pm$ 0.95 25.09$\pm$ 0.17 244.60$\pm$ 10.47 2.04$\pm$ 0.03
V3_2 75.83$\pm$ 0.83 27.14$\pm$ 2.16 223.59$\pm$ 6.15 1.81$\pm$ 0.06
V3_3 73.70$\pm$ 2.59 24.63$\pm$ 2.77 245.72$\pm$ 15.48 1.75$\pm$ 0.15
V3_4 52.38$\pm$ 3.81 17.61$\pm$ 0.82 316.57$\pm$ 4.58 2.46$\pm$ 0.12
V3_5 76.67$\pm$ 0.83 31.29$\pm$ 1.91 211.91$\pm$ 5.29 1.62$\pm$ 0.04
@@
Table 14: NOLO(M) performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
@@
T10_5 80.59$\pm$ 4.12 32.92$\pm$ 1.93 202.68$\pm$ 22.39 1.47$\pm$ 0.12
T11_5 77.33$\pm$ 2.67 36.00$\pm$ 3.27 218.04$\pm$ 1.64 1.64$\pm$ 0.01
T12_5 78.82 36.62$\pm$ 0.82 211.44$\pm$ 9.89 1.55$\pm$ 0.01
T1_5 50.83$\pm$ 5.83 27.99$\pm$ 2.85 302.14$\pm$ 14.69 2.35$\pm$ 0.10
T2_5 69.05$\pm$ 1.43 25.09$\pm$ 2.83 256.49$\pm$ 1.55 1.85$\pm$ 0.13
T3_5 54.71$\pm$ 0.59 21.49$\pm$ 0.94 312.69$\pm$ 3.18 2.61$\pm$ 0.08
T4_5 73.64$\pm$ 2.73 25.06$\pm$ 1.03 245.16$\pm$ 17.39 1.78$\pm$ 0.08
T5_5 75.45$\pm$ 0.91 31.23$\pm$ 1.14 234.16$\pm$ 1.94 1.75$\pm$ 0.07
T6_5 65.83$\pm$ 2.50 24.81$\pm$ 0.36 276.04$\pm$ 19.72 1.95$\pm$ 0.01
T7_5 70.62$\pm$ 0.63 19.64$\pm$ 0.81 281.29$\pm$ 2.88 1.77$\pm$ 0.05
T8_5 59.38$\pm$ 6.87 23.84$\pm$ 2.27 287.47$\pm$ 17.78 2.38$\pm$ 0.37
T9_5 59.00$\pm$ 7.00 31.86$\pm$ 5.95 278.41$\pm$ 28.49 2.12$\pm$ 0.22
@@
V1_1 60.00$\pm$ 3.00 25.32$\pm$ 3.23 294.85$\pm$ 21.57 2.29$\pm$ 0.09
V1_2 60.00$\pm$ 1.82 24.32$\pm$ 3.48 291.72$\pm$ 6.24 2.18$\pm$ 0.09
V1_3 82.63$\pm$ 1.58 35.75$\pm$ 3.13 200.15$\pm$ 3.24 1.41$\pm$ 0.09
V1_4 62.22$\pm$ 1.11 24.95$\pm$ 0.16 288.34$\pm$ 6.13 2.19$\pm$ 0.21
V1_5 64.58$\pm$ 4.58 31.77$\pm$ 3.05 265.98$\pm$ 16.41 2.04$\pm$ 0.16
V2_1 49.09$\pm$ 1.82 14.51$\pm$ 2.21 358.64$\pm$ 2.52 2.68$\pm$ 0.06
V2_2 66.11$\pm$ 2.78 30.06$\pm$ 1.07 261.83$\pm$ 10.29 2.11$\pm$ 0.15
V2_3 77.69$\pm$ 0.77 31.33$\pm$ 1.61 232.93$\pm$ 6.27 1.49$\pm$ 0.02
V2_4 57.69$\pm$ 1.54 23.61$\pm$ 0.69 296.56$\pm$ 8.78 2.26$\pm$ 0.07
V2_5 59.57$\pm$ 4.78 23.29$\pm$ 0.57 303.90$\pm$ 19.02 2.48$\pm$ 0.30
V3_1 70.48$\pm$ 1.90 25.14$\pm$ 1.24 254.00$\pm$ 0.10 1.87$\pm$ 0.06
V3_2 71.25$\pm$ 0.42 25.59$\pm$ 0.67 244.68$\pm$ 2.23 1.88$\pm$ 0.04
V3_3 67.78$\pm$ 0.37 24.20$\pm$ 1.78 258.00$\pm$ 2.40 1.89$\pm$ 0.03
V3_4 54.29$\pm$ 0.95 13.15$\pm$ 1.33 330.42$\pm$ 8.02 2.40$\pm$ 0.00
V3_5 67.50$\pm$ 1.67 30.21$\pm$ 4.40 246.60$\pm$ 8.73 2.05$\pm$ 0.04
@@
9.2 Habitat Evaluation Performance
Also, we examined the impact of NOLO compared to LMM baselines and visual navigation baselines on two key metrics for vision-based navigation tasks: Success Rate (SR) and Success weighted by normalized inverse Path Length (SPL), as illustrated in Fig. 11. The performances of NOLO, random policy, GPT-4o, Video-LLaVA, AVDC, VGM, ZSON, NOLO-C, NOLO-T, NOLO(M) in Habitat testing scenes are reported in Tabs. 15, 16, 17, 18, 19, 20, 21, 22, 23 and 24 respectively.
(a) Average SR
(b) Average SPL
(c) Average SR
(d) Average SPL
Figure 11: Compare NOLO with random policy, LMM baselines and visual navigation baselines in 20 Habitat testing scenes.
Table 15: NOLO performance in Habitat testing scenes.